Suppose that the sequence (sn) converges to s and sn ≤ A for every n. Show that s≤A


1
Expert's answer
2021-06-29T07:54:09-0400

Given sequence (sn) is convergent and bounded by A

i.e. sn "\leq" A for all n

(sn) is convergent therefore by definition -

("\lim_{n \to \infty}" sn = L ,if for every number  "\epsilon" > 0.there is an integer N such that "\mid" sn - L "\mid" < "\epsilon" whenever n > N)

since the sequence is bounded hence every term is less than A

and by definition of limit after n > N, "\mid" sn - L "\mid" < "\epsilon" (i.e. the difference between sn and L is negligible OR sn is almost equal to L)

therefore,

"\mid" L "\mid" "\leq" A